OpenStudy (anonymous):

proportions. To measure the distance betweens points A and B, on two islands, a man takes the following measurements on one island: AC = 5 yds, DE = 6 yds, AE = 15 yds, how far apart are the islands?

OpenStudy (anonymous):

If anyone can please check this, I did it two ways with two answers: x/15 = 6/5 5x = 75 x = 15 yds (this was the first way) 5/6 = x/15 6x = 75 x = 12.5 yds ( second way) not sure if I'm lining up my ratios properly???)

OpenStudy (amistre64):

are we to assume a flat planet?

OpenStudy (anonymous):

looks like it in the book... an island with water spaced to another island. its a right triangle from one island to the other... a line creates a right angle bisector at the end of the first island is the 5 yds

OpenStudy (amistre64):

what is DE? is that a typo?

OpenStudy (amistre64):

how can i make this look right?

OpenStudy (anonymous):

6 yards is DE thats the right angle of the big triangle. on the first island as well ... No it is a right triangle

OpenStudy (anonymous):

.D .C 6 5 .B .E .A Line EA is 15 sorry for the cheap drawing draw a line to the dots

OpenStudy (anonymous):

yes and then at the bottom tip there is a straight line CA that is 5 yds

OpenStudy (amistre64):

i can work with that

OpenStudy (anonymous):

the triangle is also a little more narrow

OpenStudy (amistre64):

like this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

amistre u rekn u cud take a quick look at my q just under this one when ur done?

OpenStudy (anonymous):

top of the triangles D the right angle is E and the tip is B

OpenStudy (anonymous):

that line in the middle is CA

OpenStudy (amistre64):

tanB = tanB 6 5 ------ = ---- ; cross multiply now 15 +x x 6x = 75+5x x = 75

OpenStudy (amistre64):

the islands are 75 apart then

OpenStudy (anonymous):

Thank you I was off!